# Singularity of the loops within a cable-graph loop-soup conditioned by its occupation time

### v1

@@ 1. Introduction
It is a rather natural question to investigate the conditional law of $(\beta_{t},t\leq T)$ when one conditions on its occupation time profile at a positive time $T$. This is the so-called Brownian burglar that was introduced and studied by Warren and Yor in [17] – the case where $\beta$ is a Brownian excursion has then been studied by Aldous in [1] (and this case will be closely related to the setups that we will be studying). In other words, one conditions the Brownian path $\beta$ by where/how it “spreads” its time until $T$. We can note that any almost sure properties of the Brownian paths (like abouts its quadratic variation, the modulus of continuity etc) will necessarily still hold for this conditioned process (indeed, an event that holds almost surely will almost surely hold almost surely for the conditional law), so that the trajectory of a Brownian burglar looks a priori similar to that of an usual Brownian path. However, as we will see in the present paper, if $T$ is the first time at which the local time at the origin of $\beta$ reaches some positive value $l$, then it is possible to argue that for all $\varepsilon>0$, the conditional law of $(\beta_{t},t\leq\varepsilon)$ given $\mu_{T}$ is singular with respect to that of an unconditioned Brownian motion up to time $\varepsilon$ (this will be Proposition 5).
@@
Our main purpose in the present paper is in fact to study this type of question in the closely related set-up of Brownian loop-soups in transient cable-graphs that we now briefly recall. A cable-graph $G$ can be viewed as the metric space obtained when one first considers a discrete graph, then associates to each edge $e$ joining two sites $x$ and $y$ in $G$ a segment $I_{e}$ isomorphic to some compact interval of $\mathbb{R}$, where the endpoints of the edges correspond to $x$ and $y$. The cable-graph is then the union of these “physical” edges $I_{e}$. On this graph, it is then possible to define Brownian motion (loosely speaking, when it is in the middle of an edge, it moves like a one-dimensional Brownian motion, and when it is at a vertex, it chooses for each excursion away from $x$ uniformly at random in which edge adjacent to $x$ to start). The trace on the sites of the graph of such a Brownian motion can be viewed as a random walk of $G$. One can then also define a natural measure on Brownian loops on such a cable-graph, and then consider a Poisson point process of Brownian loops with intensity given by some constant $c$ times this Brownian loop measure (see [8] for background). When the graph $G$ is transient, this is a Poisson point process of (unrooted) Brownian loops where each given edge is almost surely visited by infinitely many loops, but only by finitely many ones with diameter greater than any fixed $\epsilon$. Each of these Brownian loops in the loop-soup has a continuous occupation time density, and when one sums these densities over all Brownian loops in the loop-soup, one obtains the occupation time density of the entire Brownian loop-soup. This is a random function $\Lambda_{c}$ that can be viewed as the cable-graph generalization of the squared Bessel processes of dimension $c$. In particular, when $c=1$, the law of $\Lambda_{c}$ is that of the square of a Gaussian Free Field on this cable-graph (recall that the GFF on the cable-graph is a continuous function from the cable-graph into $\mathbb{R}$ that can be viewed as the natural way to extend Brownian bridges on each interval $I_{e}$ to the whole cable-graph).
@@ Acknowledgements
I thank Wendelin Werner for help and advice during the preparation of this note. This research has been funded by a grant from the Royal Society.

### v2

@@ 1. Introduction
It is a rather natural question to investigate the conditional law of $(\beta_{t},t\leq T)$ when one conditions on its occupation time profile at a positive time $T$. This is the so-called Brownian burglar that was introduced and studied by Warren and Yor in [19] – the case where $\beta$ is a Brownian excursion has then been studied by Aldous in [2]. More recently, Aïdékon, Hu and Shi [1] did use another approach based on the joint law of all local times profiles (a little bit in the spirit of the construction of the “true self-repelling motion” in [18]) to construct the process (which will be rather closely related to the setups that we will be studying). In other words, one conditions the Brownian path $\beta$ by where/how it “spreads” its time until $T$. We can note that any almost sure properties of the Brownian paths (like abouts its quadratic variation, the modulus of continuity etc) will necessarily still hold for this conditioned process (indeed, an event that holds almost surely will almost surely hold almost surely for the conditional law), so that the trajectory of a Brownian burglar looks a priori similar to that of an usual Brownian path. However, as we will see in the present paper, if $T$ is the first time at which the local time at the origin of $\beta$ reaches some positive value $l$, then it is possible to argue that for all $\varepsilon>0$, the conditional law of $(\beta_{t},t\leq\varepsilon)$ given $\mu_{T}$ is singular with respect to that of an unconditioned Brownian motion up to time $\varepsilon$ (this will be Proposition 5).
@@
Our main purpose in the present paper is in fact to study this type of question in the closely related set-up of Brownian loop-soups in transient cable-graphs that we now briefly recall. A cable-graph $G$ can be viewed as the metric space obtained when one first considers a discrete graph, then associates to each edge $e$ joining two sites $x$ and $y$ in $G$ a segment $I_{e}$ isomorphic to some compact interval of $\mathbb{R}$, where the endpoints of the edges correspond to $x$ and $y$. The cable-graph is then the union of these “physical” edges $I_{e}$. On this graph, it is then possible to define Brownian motion (loosely speaking, when it is in the middle of an edge, it moves like a one-dimensional Brownian motion, and when it is at a vertex, it chooses for each excursion away from $x$ uniformly at random in which edge adjacent to $x$ to start). The trace on the sites of the graph of such a Brownian motion can be viewed as a random walk of $G$. One can then also define a natural measure on Brownian loops on such a cable-graph, and then consider a Poisson point process of Brownian loops with intensity given by some constant $c$ times this Brownian loop measure (see [9] for background). When the graph $G$ is transient, this is a Poisson point process of (unrooted) Brownian loops where each given edge is almost surely visited by infinitely many loops, but only by finitely many ones with diameter greater than any fixed $\varepsilon$. Each of these Brownian loops in the loop-soup has a continuous occupation time density, and when one sums these densities over all Brownian loops in the loop-soup, one obtains the occupation time density of the entire Brownian loop-soup. This is a random function $\Lambda_{c}$ that can be viewed as the cable-graph generalization of the squared Bessel processes of dimension $c$. In particular, when $c=1$, the law of $\Lambda_{c}$ is that of the square of a Gaussian Free Field on this cable-graph (recall that the GFF on the cable-graph is a continuous function from the cable-graph into $\mathbb{R}$ that can be viewed as the natural way to extend Brownian bridges on each interval $I_{e}$ to the whole cable-graph).
@@ Acknowledgements
I thank Wendelin Werner for help and advice during the preparation of this note. I also thank Elie Aïdékon for his comments on the first version of this paper. This research has been funded by a grant from the Royal Society.